# A Multi-Modal Explainability Approach for Human-Aware Robots in Multi-Party Conversation

### v1

@@ Acknowledgments
To merge the two data streams, we also tested some additional architectural concepts. These methods however, neither surpassed the baseline accuracy significantly, nor brought clear explanation to the user.
Additional method 1
@@
In this part, we are going to transform the two modalities of pose and image vector. We base it on the mechanism of multi-dimensional attention proposed in Shen et al., (2017), where the authors propose an attention scoring not only to weight individual value vectors but to give each of their elements a corresponding weight as well. For a given time frame $t$, using this procedure we are able to merge two modalities (outputs of the face and pose networks, $\boldsymbol{f}_{t}\in\mathbb{R}^{d_{face}}$ and $\boldsymbol{p}_{t}\in\mathbb{R}^{d_{pose}}$ respectively) into a single representation, which is then fed to the final, recurrent network. Using this method, the modalities are combined while taking into account their cross-relations. Our divergence from the traditional approach is that we do not combine multiple value vectors to a single representation, but we rather compute one value vector at a time and feed it to the RNN. The exact computation is following:
$\boldsymbol{q}_{t}=\mathbf{W}^{Q}\boldsymbol{p}_{t},\quad\boldsymbol{k}_{t}=\mathbf{W}^{K}\boldsymbol{f}_{t},\quad\boldsymbol{v}_{t}=\mathbf{W}^{V}\boldsymbol{f}_{t},$ (C.1)
@@
where the matrices, $\mathbf{W}^{Q}\in\mathbb{R}^{d_{q}\times d_{pose}}$, $\mathbf{W}^{K}\in\mathbb{R}^{d_{k}\times d_{face}}$ and $\mathbf{W}^{V}\in\mathbb{R}^{d_{v}\times d_{face}}$ are trainable parameters representing transformation of the corresponding vectors to query, key and value.
@@
$\boldsymbol{e}_{t}=\boldsymbol{W}^{T}_{D}\times act(\boldsymbol{W}_{1}\times\boldsymbol{q}_{t}+\boldsymbol{W}_{2}\times\boldsymbol{k}_{t}+\boldsymbol{b}),$ (C.2)
where $\boldsymbol{W}_{1}\in\mathbb{R}^{d_{inner}\times d_{q}}$, $\boldsymbol{W}_{2}\in\mathbb{R}^{d_{inner}\times d_{k}}$, $\boldsymbol{W}_{D}\in\mathbb{R}^{d_{inner}\times d_{v}}$ and $\boldsymbol{b}\in\mathbb{R}^{d_{inner}}$ are trainable parameters.
@@
Additional method 2
In this method we leverage the idea of general attention procedure (Brauwers and Frasincar,, 2023). After the input is processed, we end up with a representation of pose and face, $\boldsymbol{p_{t}}$ and $\boldsymbol{f_{t}}$, respectively. However in this scenario, $\boldsymbol{f_{t}}\in\mathbb{R}^{d_{pose}}$ is a vector but $\boldsymbol{p_{t}}\in\mathbb{R}^{n_{p}\times d_{embed}}$ is a series of vectors, each corresponding to a channel aggregation for a single pixel value after the convolutions.
@@
$\boldsymbol{q}_{t}=\mathbf{W}^{Q}\boldsymbol{p}_{t},\quad\boldsymbol{k}_{t,i}=\mathbf{W}^{K}\boldsymbol{f}_{t,i},\quad\boldsymbol{v}_{t,i}=\mathbf{W}^{V}\boldsymbol{f}_{t,i},$ (C.3)
where the matrices, $\mathbf{W}^{Q}\in\mathbb{R}^{d_{q}\times d_{pose}}$, $\mathbf{W}^{K}\in\mathbb{R}^{d_{k}\times d_{face}}$ and $\mathbf{W}^{V}\in\mathbb{R}^{d_{v}\times d_{face}}$ are trainable parameters representing transformation of the corresponding vectors to query, key and value. A scoring function is applied to return the corresponding weight to each of the components of the face representation:
$\boldsymbol{e}_{t,i}=\boldsymbol{w}^{T}_{D}\times act(\boldsymbol{W}_{1}\times\boldsymbol{q}_{t}+\boldsymbol{W}_{2}\times\boldsymbol{k}_{t,i}+\boldsymbol{b}),$ (C.4)
where $\boldsymbol{W}_{1}\in\mathbb{R}^{d_{inner}\times d_{q}}$, $\boldsymbol{W}_{2}\in\mathbb{R}^{d_{inner}\times d_{k}}$, $\boldsymbol{W}_{D}\in\mathbb{R}^{d_{inner}}$ and $\boldsymbol{b}\in\mathbb{R}^{d_{inner}}$ are trainable parameters. This is followed by an alignment step, where we normalize the contribution of each component of the face representation.
$a_{t,i}=\rm{softmax}(e_{t,i};\boldsymbol{e}).$ (C.5)
@@
$\boldsymbol{r_{t}}=\sum_{i=1}^{n_{p}}a_{t,i}.\boldsymbol{v_{t,i}}.$ (C.6)
Appendix D
Prompts for the LLM agent
Shopping Mall Context: “You are a service robot named iCub, working as an assistant in shopping mall. You help customers who need information about shops inside the mall. You give informations based on customers needs and preferences. Be friendly and keep answers very short and concise!.”
Domestic Assistant Context: “You are a domestic assistant robot named iCub. You can do several tasks, like preparing drinks and food, and your role is to help accomplish this task when required. Be friendly and keep answers very short and concise!.”
Restaurant Context: “You are a service robot named iCub, working as a waiter for a restaurant. You help customers who need to order their food and drinks. Be friendly and keep answers very short and concise!.”
Appendix E
Table E.1: Results from the correlation matrix of users’ Explainability evaluation related to the embodied modality.
@@
Usefulness Spearman’s $\rho$ 0.611 **
df 19
p-value 0.003
Intrusiveness Spearman’s $\rho$0.223 -0.644 **
df 19 19
p-value 0.332 0.002
Note. * p < 0.05, ** p < 0.01, *** p < 0.001
Table E.2: Results from the correlation matrix of users’ Explainability evaluation related to the verbal modality.
@@
Usefulness Spearman’s $\rho$ 0.479 *
df 19
p-value 0.028
Intrusiveness Spearman’s $\rho$0.117 -0.332
df 19 19
p-value 0.614 0.142
Note. * p < 0.05, ** p < 0.01, *** p < 0.001
Table E.3: Results from the correlation matrix of users’ Explainability evaluation related to the visual/attentional modality.

### v2

@@ Acknowledgments
To merge the two data streams, we also tested some additional architectural concepts. However, these methods neither significantly surpassed the baseline accuracy nor brought clear explanations to the user.
Modality merge using multi-dimensional attention
@@
In this part, we are going to transform the two modalities of pose and image vector. We base it on the mechanism of multi-dimensional attention proposed in Shen et al., (2017), where the authors propose an attention scoring not only to weight individual value vectors but to give each of their elements a corresponding weight as well. For a given time frame $t$, using this procedure we are able to merge two modalities (outputs of the face and pose networks, $\boldsymbol{f}_{t}\in\mathbb{R}^{d_{\text{face}}}$ and $\boldsymbol{p}_{t}\in\mathbb{R}^{d_{\text{pose}}}$ respectively) into a single representation, which is then fed to the final, recurrent network. Using this method, the modalities are combined while taking into account their cross-relations. Our divergence from the traditional approach is that we do not combine multiple value vectors to a single representation, but we rather compute one value vector at a time and feed it to the RNN. The exact computation is following:
$\boldsymbol{q}_{t}=\mathbf{W}^{Q}\boldsymbol{p}_{t},\quad\boldsymbol{k}_{t}=\mathbf{W}^{K}\boldsymbol{f}_{t},\quad\boldsymbol{v}_{t}=\mathbf{W}^{V}\boldsymbol{f}_{t},$ (B.1)
@@
where the matrices, $\mathbf{W}^{Q}\in\mathbb{R}^{d_{q}\times d_{\text{pose}}}$, $\mathbf{W}^{K}\in\mathbb{R}^{d_{k}\times d_{\text{face}}}$ and $\mathbf{W}^{V}\in\mathbb{R}^{d_{v}\times d_{\text{face}}}$ are trainable parameters representing transformation of the corresponding vectors to query, key and value.
@@
$\boldsymbol{e}_{t}=\boldsymbol{W}^{T}_{D}\times act(\boldsymbol{W}_{1}\times\boldsymbol{q}_{t}+\boldsymbol{W}_{2}\times\boldsymbol{k}_{t}+\boldsymbol{b}),$ (B.2)
where $\boldsymbol{W}_{1}\in\mathbb{R}^{d_{\text{inner}}\times d_{q}}$, $\boldsymbol{W}_{2}\in\mathbb{R}^{d_{\text{inner}}\times d_{k}}$, $\boldsymbol{W}_{D}\in\mathbb{R}^{d_{\text{inner}}\times d_{v}}$ and $\boldsymbol{b}\in\mathbb{R}^{d_{\text{inner}}}$ are trainable parameters.
@@
Modality merge using general attention
In this method we leverage the idea of general attention procedure (Brauwers and Frasincar,, 2023). After the input is processed, we end up with a representation of pose and face, $\boldsymbol{p_{t}}$ and $\boldsymbol{f_{t}}$, respectively. However in this scenario, $\boldsymbol{f_{t}}\in\mathbb{R}^{d_{\text{pose}}}$ is a vector but $\boldsymbol{p_{t}}\in\mathbb{R}^{n_{p}\times d_{\text{embed}}}$ is a series of vectors, each corresponding to a channel aggregation for a single pixel value after the convolutions.
@@
$\boldsymbol{q}_{t}=\mathbf{W}^{Q}\boldsymbol{p}_{t},\quad\boldsymbol{k}_{t,i}=\mathbf{W}^{K}\boldsymbol{f}_{t,i},\quad\boldsymbol{v}_{t,i}=\mathbf{W}^{V}\boldsymbol{f}_{t,i},$ (B.3)
where the matrices, $\mathbf{W}^{Q}\in\mathbb{R}^{d_{q}\times d_{\text{pose}}}$, $\mathbf{W}^{K}\in\mathbb{R}^{d_{k}\times d_{\text{face}}}$ and $\mathbf{W}^{V}\in\mathbb{R}^{d_{v}\times d_{\text{face}}}$ are trainable parameters representing transformation of the corresponding vectors to query, key and value. A scoring function is applied to return the corresponding weight to each of the components of the face representation:
$\boldsymbol{e}_{t,i}=\boldsymbol{w}^{T}_{D}\times act(\boldsymbol{W}_{1}\times\boldsymbol{q}_{t}+\boldsymbol{W}_{2}\times\boldsymbol{k}_{t,i}+\boldsymbol{b}),$ (B.4)
where $\boldsymbol{W}_{1}\in\mathbb{R}^{d_{\text{inner}}\times d_{q}}$, $\boldsymbol{W}_{2}\in\mathbb{R}^{d_{\text{inner}}\times d_{k}}$, $\boldsymbol{W}_{D}\in\mathbb{R}^{d_{\text{inner}}}$ and $\boldsymbol{b}\in\mathbb{R}^{d_{\text{inner}}}$ are trainable parameters. This is followed by an alignment step, where we normalize the contribution of each component of the face representation.
$a_{t,i}=\rm{softmax}(e_{t,i};\boldsymbol{e}).$ (B.5)
@@
$\boldsymbol{r_{t}}=\sum_{i=1}^{n_{p}}a_{t,i}.\boldsymbol{v_{t,i}}.$ (B.6)
Appendix C - Deployment of the XAE in a modular architecture: Prompts for the LLM agent
Shopping Mall Context:
“You are a service robot named iCub, working as an assistant in shopping mall. You help customers who need information about shops inside the mall. You give informations based on customers needs and preferences. Be friendly and keep answers very short and concise!.”
Domestic Assistant Context:
“You are a domestic assistant robot named iCub. You can do several tasks, like preparing drinks and food, and your role is to help accomplish this task when required. Be friendly and keep answers very short and concise!.”
Restaurant Context:
“You are a service robot named iCub, working as a waiter for a restaurant. You help customers who need to order their food and drinks. Be friendly and keep answers very short and concise!.”
LLM-based AE Predictions Context:
“You are a robot taking part in a conversation with one or more people. Given who is speaking (the Speaker) and what the speaker says (the Message), identify who is the sentence’s addressee. Choose among only two options: [Myself, the ROBOT], [Another person, NOT THE ROBOT]. Answer in the format: Addressee: [option].”
Appendix D - Online User Study: Additional Analysis
Analysis of relationships between the three explanation evaluation criteria via correlation matrices
Table D.1: Correlation matrix in the verbal modality.
@@
Usefulness Spearman’s $\rho$512 ***
df 58
p-value <.001
Intrusiveness Spearman’s $\rho$.293 * -.221
df 58 58
p-value .023 .089
Note. * p < .05, ** p < .01, *** p < .001
Table D.2: Correlation matrix in the embodied modality.
@@
Usefulness Spearman’s $\rho$770 ***
df 58
p-value <.001
Intrusiveness Spearman’s $\rho$.285 * -.262 *
df 58 58
p-value .027 .043
Note. * p < .05, ** p < .01, *** p < .001
Table D.3: Correlation matrix in the graphical modality.